OpenStudy (lilai3):

Please help! It will be greatly appreciated! Will award medal! Jerry rents snowmobiles to tourists. He charges $135 for 4 hours and $202.50 for 6 hours. What is the hourly rate Jerry charges to rent a snowmobile?

OpenStudy (mertsj):

You need to find dollars per hour. That's what hourly rate means.

OpenStudy (mertsj):

Dollars per hour means dollars over hours. So put 135 over 4 and do the DIVISION

OpenStudy (lilai3):

uh.. is it 33.75 per hour?

OpenStudy (mertsj):

Yes. Good job.

OpenStudy (mertsj):

(That's a lot isn't it?)

OpenStudy (lilai3):

yep. thanks for helping me. would you also be willing to help me on two more problems like this? it's just a check.

OpenStudy (mertsj):

yep

OpenStudy (lilai3):

The power absorbed by a solar panel varies directly with its area. If an 8 square meter panel absorbs 8,160 watts of power, how much power does a 12 square meter solar panel absorb?

OpenStudy (mertsj):

Power varies directly with area mean that power is equal to area times some other number called a constant and usually designated by k.

OpenStudy (mertsj):

So from the first sentence we can write P = kA

OpenStudy (lilai3):

what did you get?

OpenStudy (mertsj):

12240

OpenStudy (lilai3):

omg me too. one more problem?

OpenStudy (mertsj):

ok

OpenStudy (lilai3):

Mr. Malone used 40 pounds of insecticide to cover 1, 760 square feet of lawn and 60 pounds to cover an additional 2,640 square feet. How many pounds of insecticide would Mr. Malone need to cover his whole lawn of 4,480 square feet?

OpenStudy (mertsj):

101.8

OpenStudy (lilai3):

wait.. how did you get that? i got 4, 298.19....????

OpenStudy (mertsj):

1760/40=44 square feet per pound 2640/60=44 square feet per pound 4480/x=44/1 44x=4480 4480/44=101.8

OpenStudy (lilai3):

oh. thank you very much! i greatly apprecaite it. your help means a lot to me. thank you XD

OpenStudy (mertsj):

yw